ECD signal stuck at 844883.8

Hi guys,

So I was walking by one of our less frequently used 8890s and saw that all my flow and temperature zones were off and the detectors were saturated at 850k. I'm the only person who operates it and I didn't leave it like that. Checked the system log and it was fine so I started up carrier and aux gases, only problem was the front detector signal didn't move. I tried restarting the instrument, restarting openlab, turned the flow on and off for the aux gas, and pushing large volumes through both the carrier and aux lines.

Is this thing toast? It's only 14 months old.

8/23/2021 edit: I checked the exhaust flows and everything was where it should be so I heated the detectors and made a solvent injection. Back channel looked fine but the front channel never moved.

Parents
  • Pulled the electronics cover and found a loose connection that goes to the front electrometer. The black ribbon cable on top was only locked in on one side, turned the instrument on and am still getting the same reading. Is there a battery on this board somewhere that's still holding onto settings from when that cable was still loose?

                                           

Reply
  • Pulled the electronics cover and found a loose connection that goes to the front electrometer. The black ribbon cable on top was only locked in on one side, turned the instrument on and am still getting the same reading. Is there a battery on this board somewhere that's still holding onto settings from when that cable was still loose?

                                           

Children
Was this helpful?